40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 30 18:44:11 2021, Anchal, Summary, LSC, Tried fixing ETMY QPD 
    Reply  Thu Jul 1 10:34:51 2021, Paco, Anchal, Summary, LSC, ETMY QPD fixed 
Message ID: 16232     Entry time: Wed Jun 30 18:44:11 2021     Reply to this: 16233  
Author: Anchal 
Type: Summary 
Category: LSC 
Subject: Tried fixing ETMY QPD 

I worked in Yend station, trying to get the ETMY QPD to work properly. When I started, only one (quadrant #3) of the 4 quadrants were seeing any lights. By just changing the beam splitter that reflects some light off to the QPD, I was able to get some amount of light in quadrant #2. However, no amount of steering would show any light in any other quadrants.

The only reason I could think of is that the incoming beam gets partially clipped as it seems to be hitting the beam splitter near the top edge. So for this to work properly, a mirror upstream needs to be adjusted which would change the alignment of TRX photodiode. Without the light on TRX photodiode, there is no lock and there is no light. So one can't steer this beam without lossing lock.

I tried one trick, in which, I changed the YARM lock trigger to POY DC signal. I got it to work to get the lock going even when TRY was covered by a beam finder card. However, this lock was still bit finicky and would loose lock very frequently. It didn't seem worth it to potentially break the YARM locking system for ETMY QPD before running this by anyone and this late in evening. So I reset everything to how it was (except the beam splitter that reflects light to EMTY QPD. That now has equal ligth falling on quadrant #2 and #3.

The settings I temporarily changed were:

  • C1:LSC-TRIG_MTRX_7_10 changed from 0 to -1 (uses POY DC as trigger)
  • C1:LSC-TRIG_MTRX_7_13 changed from 1 to 0 (stops using TRY DC as trigger)
  • C1:LSC-YARM_TRIG_THRESH_ON changed from 0.3 to -22
  • C1:LSC-YARM_TRIG_THRESH_OFF changed from 0.1 to -23.6
  • C1:LSC-YARM_FM_TRIG_THRESH_ON changed from 0.5 to -22
  • C1:LSC-YARM_FM_TRIG_THRESH_OFF changed from 0.1 to -23.6

All these were reverted back to there previous values manually at the end.

 

ELOG V3.1.3-